OpenStudy (anonymous):

Where should the point P be chosen on line segment AB so as to maximize the angle θ? (Assume a = 2 units, b = 4 units, and c = 6 units. Round your answer to two decimal places.)

OpenStudy (jmark):

Given a = 2 units, b = 4 units, and c = 6 units Let x = measure of PB. This makes PA = 4 - x. Let C = other point of the triangle with P and B. Let y = <PCB. Let D = other point of the triangle with P and A. Let z = <PDA. Notice that theta = y + z. tan y = x/a = x/2 y =arctan(x/2) tan z = (4 - x)/c = (4 - x)/6 z = arctan(4 - x)/6 Therefore, theta is a function of x, namely f(x) = arctan(x/2) + arctan(4 - x)/6 f(x)'s maximum occurs where f '(x) = 0. f '(x) = 1/2(1/(1 + (x/2)^2) - 1/6(1/(1 + ((4 – x)/6)^2)) f'(x)=2/(x^2+4)-3/(x^2-8x+52) f '(x) = 0 when =>2/(x^2+4)-3/(x^2-8x+52)=0 x = 2(-4-√39) x=-20.49 or x = 2(√39-4) x=4.49 A negative length makes no sense so discard x = -20.49, leaving just x = 4.49 Know more about Decimal Division http://goo.gl/z99j9s